If the length of a rectangle is 3 times its width, and if its area is 108, what is its perimeter?

Answer by mananth(16946)

width=x
length = 3x
3x*x=108
3x^2=108
x^2=36
x=6 the width
Length = 18
..
Perimeter = 2*(L+W)
Perimeter =2*(6+18)
Perimeter = 48 units